# SonicID: User Identification on Smart Glasses with Acoustic Sensing

### v1

@@ 3.2. Signal Transmission and Reception
In SonicID, we choose signals that sweep within a certain frequency range as transmitted signals. Considering SonicID is deployed on smart glasses which is quite close to users when they put it on, we decide to sweep the signals in the ultrasonic frequency bands, i.e. over 18 kHz, to alleviate the impact of the audible sound on users’ daily activities. In order to obtain richer information from users’ face, we put one pair of speaker and microphone on each side of the glasses. The speakers on two sides are designed to transmit encoded signals in different ultrasonic frequency ranges to avoid interference. We pick 18-21 kHz for the speaker on the right side and 21.5-24.5 kHz for the speaker on the left so that they are inaudible to most people and compatible with the audio interfaces on most commodity devices. A sampling rate of 50 kHz is selected to support these two frequency ranges and the signals sweep from the lower frequency boundary (18/21.5 kHz) to the higher frequency boundary (21/24.5 kHz) every 12 ms. Therefore, each sweep contains N = 600 samples ($50kHz\times 12ms$). Fig. 1 (a) demonstrates one of the transmitted signals that sweeps from 18-21 kHz. These signals are commonly used and named as Frequency Modulated Continuous Waves (FMCW) in the field.
@@ 3.3. Echo Profile Calculation
The two microphones capture two channels of audio on two sides of the glasses. An example of the received signal in one channel (18-21 kHz) is shown in Fig. 1 (b). After receiving the signals, we first apply two Butterworth band-pass filters, which are from 18-21 kHz and 21.5-24.5 kHz, on the signals separately to filter out the noises that are outside the frequency range of our interest. By applying two band-pass filters on two channels of signals, we get four channels of filtered signals, two of which represent the signals that travel on the same side of the face while the other two represent the signals that travel across the face from one side to the other side of the face. The acoustic patterns associated to the user’s face that we would like to obtain are called Echo Profiles, which have been demonstrated in several previous papers (Wang et al., 2018; Li et al., 2022; Mahmud et al., 2023; Appear], 2024). According to prior work, we continuously calculate the cross-correlation between the received signals and the transmitted signals using Eq. 1.
@@
We then reshape the one-dimensional $C(n)$ by the frame size N = 600 samples to form a two-dimensional array. This array of patterns is defined as Echo Profile, which includes the biometric information of the face of the user who is wearing the device. Fig. 1 (d) shows an example of the echo profile that is properly reshaped from Fig. 1 (c). In echo profiles, the y-axis is the distance, which shows the strength of the signals that is reflected from different distances away from the SonicID system on glasses. Because different users have different facial shape and contours, the echo profile shows the facial scanning result of each user and is distinctive for each user. Considering the average length of people’s faces, we crop 70 samples of echo profiles from -15 samples to 55 samples as the acoustic pattern to authenticate users instead of using the full range of it, which scans the area at a distance of 0 cm to 18.7 cm from the system ($55~{}samples\div 50kHz\times 340m/s\div 2$). We divide the distance by 2 because the signals travel round-way from the system to the face and back to the system. 18.7 cm is the maximum one-way distance from the system that we would like to scan. As stated above, the negative samples are from previous frames and theoretically do not contain much useful information. However, we still include 15 samples from the negative side because our transmitted signals are not infinite in the frequency domain and thus the acoustic patterns diffuse to the negative side to some extent when the cross-correlation is calculated. The cropped echo profiles are fed into a deep learning model to authenticate users, which is introduced in the subsection below.
@@ 3.4. Machine Learning Model for Authentication
Though the acoustic data is originally one-dimensional, we reshape the echo profiles to make them two-dimensional images. In order to extract features that are unique to each user from these two-dimensional echo profiles, we decide to adopt a deep learning model based on the ResNet-18 architecture because Convolutional Neural Networks (CNN) are known for being good at distinguishing different classes. Hence, we construct the model using the ResNet-18 architecture as the encoder to extract features from echo profiles and using a fully-connected layer as the decoder. The model serves as a binary classifier which classifies the real user as positive and the attackers as negative. While training the model for a new user, the positive instances come from the data that this user provides when he/she registers on the device and the negative instances come from an existing dataset which was pre-collected from different people.
The input fed into the model is the reshaped and cropped echo profiles. We take 10 600-sample frames as one instance, which corresponds to 0.12 s in time ($10\times 600~{}samples\div 50kHz$). Therefore, the size of each instance as input is $4\times 10\times 70$, where 4 represents the 4 channels of audio data and 70 represents the number of the cropped vertical samples. During testing, the positive instances are still provided by the real user while the negative instances come from the unknown attackers that have not been seen by the model.
In the CNN model, we use an Adam optimizer, a cosine scheduler and set the initial learning rate to be 0.0002. The model is trained to minimize the cross-entropy loss for 50 epochs. Random vertical shift is included and random noise is added as the data augmentation methods to boost the generalizability of the model.
@@ 4.2. Form Factor
After we select the appropriate components for the SonicID system, the next step is to prototype SonicID on the glasses form factor. In order to validate the feasibility of SonicID, we decide to deploy the system on a pair of commodity glasses. Since we want to scan the user’s face to obtain their biometric information, we would like to point the speakers and microphones downwards, emitting signals directly towards the user’s face. There are only limited space on glasses where we can place the sensors, facing downwards, as shown in Fig. 2 (b): (1) the bridge of the glass frame; (2) the bottom of the glass frame; (3) the hinges of the glasses. If the sensors are placed at position (1) or (2), there is a chance that the sensors may touch the user’s face or block the view of the user because these two positions are right in front of the user’s face. Therefore, we determine to put the sensors at the hinges of the glasses, where there is some space to instrument sensors without letting them touch the user’s face or block the user’s view. Moreover, this sensor position has been leveraged in multiple prior work to place acoustic sensors to track facial expressions (Appear], 2024) and body poses (Mahmud et al., 2023). We can potentially integrate SonicID into their systems to realize activity tracking and user authentication with one set of hardware.
As displayed in Fig. 2 (c), we put one pair of speaker and microphone on each side of the glasses to obtain richer information and the speaker and microphone are close to each other to better capture the reflected signals. The micro-controller Teensy 4.1 is also fixed to the left leg of the glasses with hot glue and tapes to make the entire system wearable. The FPC cables connecting Teensy 4.1 and the speakers and microphones are properly restrained and aligned to the glass frame. The final prototype looks similar to a pair of commercial smart glasses and is comfortable to wear, as shown in Fig. 2 (d).
@@ 5. User Study
To evaluate the performance of SonicID, we designed and conducted a user study across several days. The goal of the study is to validate SonicID’s capability of authenticating users and protecting the system from attacks.
@@ 5.1. Study Design
After the prototype was properly worn, the data collection process started. An instruction video was played on the laptop. The participants were first instructed to stay still for 10 seconds when the system scanned their face continuously and then asked to remount the device within 10 seconds. Remounting meant taking off the prototype and putting it back on. This procedure was designed to simulate real-life wearing experience since users of wearable devices frequently remount the device everyday. This added variance to the collected data. After the prototype was remounted, the participants stayed still for another 10 seconds and repeated the process above. The data collection lasted for 12 minutes for each participant where 36 10-second remounting sessions of data was collected.
The participants were asked to come back on two other days after the first day and collected data two more times with the same procedures because we would like to evaluate the stability of SonicID across different days. After the participants were done with the study on all three days, they filled out a questionnaire to provide their demographic information and feedback on using the prototype and the system.
@@ 5.2. Participants
The user study was approved by the Institutional Review Board (IRB) of the home institution of the researchers. We successfully recruited 24 participants (13 females, 9 males and 2 non-binaries) with an average age of 23.8 years old, ranging from 20 years old to 30 years old. For each participant, we collected 18 minutes of data, which were divided in 108 10-second remounting sessions across three days. Hence, 432 minutes of data was collected in total for this user study. Note that the data P11 collected on Day 3 was lost due to the hardware issue so we asked the participant to come back again and redo the study on Day 3.
@@ 6.1.1. Cross-session Performance
In the user study, we collected 36 sessions (6 minutes) of data from each participant on each day. We first evaluated SonicID’s performance across different remounting sessions on the same day. Out of the 36 sessions we collected on Day 1, we considered the first 6 sessions (1 minute) of data as the practice sessions and dumped them while we evaluated the system. Among the remaining 30 sessions, we used the first 24 sessions (4 minutes) of data as positive samples to train the deep learning model described in Sec. 3.4. We would like to scan users’ face while they are staying still. However, they can move and/or blink unconsciously during the study, which might have an impact on the scanning results. Therefore, we divided each 10-second session into multiple instances of 0.12 s, as discussed in Sec. 3.4, and picked top five instances with the smallest movements by comparing the energy change within these instances. Five 0.12-second instances from each session were actually used for training the model. The negative samples used to train the model came from a dataset we collected prior to the user study. To collect this dataset, we ran a pilot study of the same procedures as the formal user study with 17 people, including 7 researchers and 10 people from their networks. Among these people, 7 of them collected data on three days, 2 collected data on two days and the remaining 8 collected data on one day. Each person contributed 5 minutes of data of staying still on each day. All these data in the dataset were used for training the model as negative samples. The same process was conducted to select the most static instances from all the sessions while training.
Figure 3. True Positive Rate for all Participants across Three Days.
Figure 4. False Positive Rate for all Participants across Three Days.
Figure 5. Balanced Accuracy for all Participants across Three Days.
@@
After the model was properly trained for each participant in the user study, the remaining 6 sessions (1 minute) of data from this participant was adopted as positive samples to test the model. The negative samples for testing came from the other 23 participants in the user study. Each of these 23 participants contributed 5 minutes of data to test the model. In this way, these "attackers" during evaluation were not seen by the model in the training process. Static instances were also selected from all sessions for testing. The model made binary classifications to authenticate users and computed the three metrics introduced in Sec. 3.5 to evaluate the results. We trained an individual model for each participant and the average TPR, FPR and BAC across 24 participants is 96.5% (std=9.0%), 4.1% (std=4.6%) and 96.2% (std=5.6%). The detailed metrics for every participant are shown in Fig. 3, Fig 4 and Fig. 5 respectively.
@@
The average FPR of 4.1% indicates that the system rejects false users with a success rate of over 95%. It demonstrates the ability of SonicID to protect the smart glasses from being attacked by bad actors. The average TPR of 96.5% means that the system authenticates the real users successfully with a rate of over 95%. It shows the ability of SonicID to recognize true users of the smart glasses. BAC gives a balanced evaluation of the two metrics above.
@@ 6.1.2. Cross-day Performance
The results above validated the performance of SonicID across different remounting sessions on the same day. Then we evaluated the performance across different days, which was meant to validate the stability of SonicID. On Day 2 and Day 3, we also collected 6 minutes of data from each participant using the same procedures as Day 1. To be consistent with Day 1, we also dropped the first 1 minute of data from every participant. Then the static instances picked from the remaining 5 minutes of data were used for testing the cross-day performance of SonicID. We did not retrain the models and directly used the models trained on Day 1. Each participant provided his/her own 5 minutes of data as positive samples to test the model for TPR on Day 2 and Day 3 separately. The negative samples still came from the other 23 participants in the user study to test FPR.
@@
The results showed that the average TPR, FPR and BAC for Day 2 across 24 participants are 91.3% (std=11.9%), 4.5% (std=4.8%) and 93.4% (std=6.2%) while those for Day 3 are 80.7% (std=18.3%), 4.2% (std=5.1%) and 88.3% (std=10.1%). The individual study results for each participant are displayed in Fig. 3, Fig 4 and Fig. 5 respectively. We ran a one-way repeated measures ANOVA test among TPR on three days for all 24 participants and identified a statistically significant difference ($F(2,69)=7.99,p=0.001<0.05$). Specifically, we conducted a repeated measures t-test between TPR on Day 1 and Day 2, and we did not find a quite statistically significant difference ($t(23)=1.75,p=0.09>0.05$). While the same repeated measures t-test was conducted between TPR on Day 1 and Day 3, a statistically significant difference was found ($t(23)=4.00,p=0.0006<0.05$). This indicates that TPR remains consistent on Day 2 but decreases on Day 3, compared to that on Day 1. However, the average TPR is still over 80% even on Day 3. Under the cases when the real users fail to authenticate themselves into the smart glasses, they can adjust the glasses position or remount the device once, and they will likely log into the system successfully. The decrease of TPR can be attributed to the lack of positive training data from the user. One potential solution to improving TPR is to continuously collect the biomentric information of users’ face while they are using the device and update the model accordingly to increase the success rate of authentication. We plan to explore this more in the future.
@@
Similarly, we also ran a one-way repeated measures ANOVA test among FPR on three days for all 24 participants and did not find a statistically significant difference ($F(2,69)=0.39,p=0.68>0.05$). This validates the stability of SonicID to protect the system from attackers across different days. In such authentication systems, it is usually more crucial to keep FPR lower than keeping TPR higher in case of a tradeoff since we generally do not want other people to have easy access to our personal devices.
@@
In Fig. 3 and Fig. 4, we can see that P20 has the lowest TPR, especially on Day 3 while P12 has the highest FPR. By checking the videos of the participants recorded during the study for reference, we figured that P12 and P20 have long hair and bangs which made it harder for them to put the glasses back to the same position every time after remounting. This negatively affected the performance of the system on them. In Sec. 7.4, we discussed several potential methods to improve the system performance and stability in future.

### v3

@@ 3.2. Signal Transmission and Reception
In SonicID, we choose signals that sweep within a certain frequency range as transmitted signals. Considering SonicID is deployed on smart glasses which is quite close to users when they put it on, we decide to sweep the signals in the ultrasonic frequency bands, i.e. over 18 kHz, to alleviate the impact of the audible sound on users’ daily activities. In order to obtain richer information from users’ face, we put one pair of speaker and microphone on each side of the glasses. The speakers on two sides are designed to transmit encoded signals in different ultrasonic frequency ranges to avoid interference. We pick 18-21 kHz for the speaker on the right side and 21.5-24.5 kHz for the speaker on the left so that they are inaudible to most people and compatible with the audio interfaces on most commodity devices. According to the Nyquist sampling theorem, the sampling rate of the system should be at least twice the highest frequency it supports to avoid alias. Hence, the lower bound of the sampling rate for our system is $2\times 24.5kHz=49kHz$. In practice, the sampling rate needs to be slightly larger than this lower bound to allow some room for potential frequency discrepancy. Therefore, a sampling rate of 50 kHz is selected to support these two frequency ranges. We believe that this sampling rate can be supported by the audio interfaces on most current commodity devices. A higher sampling rate such as 96 kHz or 192 kHz is not selected because we would like to keep the technical requirements of SonicID as low as possible to make it easier to be integrated into future smart glasses, as discussed in Sec. 2.3. In our design, the signals sweep from the lower frequency boundary (18/21.5 kHz) to the higher frequency boundary (21/24.5 kHz) every 12 ms. Therefore, each sweep contains N = 600 samples ($50kHz\times 12ms$). Fig. 1 (a) demonstrates one of the transmitted signals that sweeps from 18-21 kHz. These signals are commonly used and named as Frequency Modulated Continuous Waves (FMCW) in the field.
@@ 3.3. Echo Profile Calculation
The two microphones capture two channels of audio on two sides of the glasses. An example of the received signal in one channel (18-21 kHz) is shown in Fig. 1 (b). After receiving the signals, we first apply two Butterworth band-pass filters, which are from 18-21 kHz and 21.5-24.5 kHz, on the signals separately to filter out the noises that are outside the frequency range of our interest. By applying two band-pass filters on two channels of signals, we get four channels of filtered signals, two of which represent the signals that travel on the same side of the face while the other two represent the signals that travel across the face from one side to the other side of the face. The acoustic patterns associated to the user’s face that we would like to obtain are called Echo Profiles, which have been demonstrated in several previous papers (Wang et al., 2018; Li et al., 2022; Mahmud et al., 2023; Li et al., 2024a). According to prior work, we continuously calculate the cross-correlation between the received signals and the transmitted signals using Eq. 1.
@@
We then reshape the one-dimensional $C(n)$ by the frame size N = 600 samples to form a two-dimensional array. This array of patterns is defined as Echo Profile, which includes the biometric information of the face of the user who is wearing the device. Fig. 1 (d) shows an example of the echo profile that is properly reshaped from Fig. 1 (c). In echo profiles, the y-axis is the distance, which shows the strength of the signals that is reflected from different distances away from the SonicID system on glasses. Because different users have different facial shape and contours, the echo profile shows the facial scanning result of each user and is distinctive for each user. Considering the average length of people’s faces, we crop 70 samples of echo profiles from -15 samples to 55 samples as the acoustic pattern to authenticate users instead of using the full range of it, which scans the area at a distance of 0 cm to 18.7 cm from the system ($55~{}samples\div 50kHz\times 340m/s\div 2$). We divide the distance by 2 because the signals travel round-way from the system to the face and back to the system. 18.7 cm is the maximum one-way distance from the system that we would like to scan. As stated above, the negative samples are from previous frames and theoretically do not contain much useful information. However, we still include 15 samples from the negative side because our transmitted signals are not infinite in the frequency domain and thus the acoustic patterns diffuse to the negative side to some extent when the cross-correlation is calculated.
To achieve an optimal authentication performance, the system should obtain the scanned facial features of users when they are static to eliminate the impact of environmental factors. However, it is possible for the users to move and/or blink intentionally and/or unconsciously during the usage of the system, which might have a negative impact on the scanning results. Therefore, we decided to adopt multiple static short instances during which the users are mostly static to authenticate them instead of using one long instance. Based on our pilot study, we empirically chose five of the aforementioned 600-sample frames as one instance, which corresponds to 0.06 seconds in time ($5\times 600~{}samples\div 50kHz$). In our user study, we will divide one session of data into multiple 0.06-second instances to evaluate the system. These instances of cropped echo profiles are fed into a deep learning model to authenticate users, which is introduced in the subsection below.
@@ 3.4. Machine Learning Model for Authentication
Though the acoustic data is originally one-dimensional, we reshape the echo profiles to make them two-dimensional images. In order to extract features that are unique to each user from these two-dimensional echo profiles, we decide to adopt a deep learning model based on the ResNet-18 architecture because Convolutional Neural Networks (CNN) are known for being good at distinguishing different classes based on visual representations. Besides, ResNet-18 has a proper amount of parameters (around 11 million parameters), which is suitable for our dataset size and potential embedded on-chip deployment. Hence, we construct the model using the ResNet-18 architecture as a feature extractor to extract features from echo profiles and using a fully-connected layer as a classifier to distinguish samples from different classes, as shown in Fig. 2. We adopt a two-stage training method to optimize the system performance. First, we created a dataset from 21 users before the formal user study. This dataset is used to pretrain a base model, with the feature extractor and the classifier combined, to tackle a task of multi-class classification. By aiming to classify these 21 users with the highest accuracy, a powerful feature extractor is trained to extract distinguishable features from the acoustic patterns obtained by scanning each user’s face. The base model only needs to be trained once and can be used for the individual models of all new users. In the second stage, the trained feature extractor is retained while the multi-class classifier is replaced with a binary classifier. The model, after further training, classifies the real user as positive and the attackers as negative. While training the individual model for a new user, the positive instances come from the data that this user provides when he/she registers on the device and the negative instances come from the existing dataset which was pre-collected from the 21 different users. This two-stage training technique enhances the feature extractor to focus more on the delicate differences across various users, compared with directly training a binary classifier.
Figure 2. Machine Learning Pipeline for SonicID.
As introduced in Sec. 3.3, the input fed into the model is the reshaped and cropped echo profiles. We take 5 600-sample frames as one instance. Therefore, the size of each instance as input is 4×4$\times$4 × 5 ×$\times 70$, where 4 represents the 4 channels of audio data and 70 represents the number of the cropped vertical samples. During testing, the positive instances are still provided by the real user while the negative instances come from the unknown attackers that have not been seen by the model during training.
In the CNN model, we use an Adam optimizer, a cosine scheduler and set the initial learning rate to be 0.0002. The model is trained to minimize the cross-entropy loss for 20 epochs for the base model and for another 10 epochs for the individual models. Random vertical shift is included and random noise is added as the data augmentation methods while the model is trained to boost its generalizability. Both of these data augmentation methods are performed at the instance level. For each instance, we implement random vertical shift by first cropping 110 vertical samples and randomly selecting 70 consecutive samples out of these 110 samples to mitigate the impact of different wearing positions of the glasses. Theoretically, the augmented dataset is 40 times ($110-70$) larger than the original dataset after random vertical shifts if the model is trained for enough epochs. Random noise is introduced by multiplying each data point of the input instance with a random factor from 0.9 to 1.1 with a probability of 80%. Moreover, z-score normalization is applied on each instance with a probability of 50%. This method improves the generalizability of the model across different wearing sessions and different days while guaranteeing that there are enough original training samples in the meantime. With the existence of random noise, the augmented dataset is at least two times larger (normalized/non-normalized).
@@ 4.2. Form Factor
After we select the appropriate components for the SonicID system, the next step is to prototype SonicID on the glasses form factor. In order to validate the feasibility of SonicID, we decide to deploy the system on a pair of commodity glasses. Since we want to scan the user’s face to obtain their biometric information, we would like to point the speakers and microphones downwards, emitting signals directly towards the user’s face. There are only limited space on glasses where we can place the sensors, facing downwards, as shown in Fig. 3 (b): (1) the bridge of the glass frame; (2) the bottom of the glass frame; (3) the hinges of the glasses. If the sensors are placed at position (1) or (2), there is a chance that the sensors may touch the user’s face or block the view of the user because these two positions are right in front of the user’s face. Therefore, we determine to put the sensors at the hinges of the glasses, where there is some space to instrument sensors without letting them touch the user’s face or block the user’s view. Moreover, this sensor position has been leveraged in multiple prior work to place acoustic sensors to track facial expressions (Li et al., 2024a) and body poses (Mahmud et al., 2023). We can potentially integrate SonicID into their systems to realize activity tracking and user authentication with one set of hardware.
As displayed in Fig. 3 (c), we put one pair of speaker and microphone on each side of the glasses to obtain richer information and the speaker and microphone are close to each other to better capture the reflected signals. The micro-controller Teensy 4.1 is also fixed to the left leg of the glasses with hot glue and tapes to make the entire system wearable. The FPC cables connecting Teensy 4.1 and the speakers and microphones are properly restrained and aligned to the glass frame. The final prototype looks similar to a pair of commercial smart glasses and is comfortable to wear, as shown in Fig. 3 (d). We used a scale to measure the weight of the prototype. The prototype itself is weighed to be 48.5 grams in total while the pair of commercial glasses we purchased are 27.8 grams and the MCU and sensors on it are 20.7 grams. If a Li-Po battery of 150 mAh capacity is included into the system, the total weight of the prototype will be 52.3 grams, which we assume is comfortable to wear for general users.
@@ 5. User Study
To evaluate the performance of SonicID, we designed and conducted a user study across several days and various scenarios. The goal of the study is to validate SonicID’s capability of authenticating users and protecting the system from attacks.
@@ 5.1. Study Design
After the prototype was properly worn, the data collection process started. An instruction video was played on the laptop. The participants were first instructed to sit still for 10 seconds when the system scanned their face continuously and then asked to remount the device within 10 seconds. Remounting meant taking off the prototype and putting it back on. This procedure was designed to simulate real-life wearing experience since users of wearable devices frequently remount the device everyday. This added variance to the collected data. After the prototype was remounted, the participants stayed still for another 10 seconds and repeated the process above. The data collection lasted for 12 minutes for each participant where 36 10-second remounting sessions of data was collected. For a selected number of participants, after the data collection process above, we asked them to repeat the aforementioned process for 2 minutes respectively when they were talking, standing and lying down. The participants first kept talking while the system scanned their face. Then they stood up and stood still while the system was in operation. Lastly, they lied down on a bean bag to evaluate the system. This design aimed to evaluate the performance of SonicID in various application scenarios.
All the participants were asked to come back on two other days after the first day and collected data two more times with the same procedures because we would like to evaluate the stability of SonicID across different days. After the participants were done with the study on all three days, they filled out a questionnaire to provide their demographic information and feedback on using the prototype and the system.
@@ 5.2. Participants
The user study was approved by the Institutional Review Board (IRB) of the home institution of the researchers. We successfully recruited 40 participants (21 females, 17 males and 2 non-binaries) with an average age of 22.6 years old, ranging from 18 years old to 30 years old. During the recruitment of participants, we did not set any criteria to exclude any specific participant except that they should be at least 18 years old. The set of participants is mostly gender balanced and covers diverse ethnic groups. Furthermore, SonicID utilizes a sound-based technique to authenticate users so the skin colors of participants should not make a large difference in terms of system performance in theory, compared with camera-based methods. As a result, we anticipate that SonicID is capable of being scaled to a much larger group of users with different backgrounds.
For each participant, we collected 18 minutes of data while they were sitting, which were divided in 108 10-second remounting sessions across three days. For the last 16 participants, we collected another 9 minutes of data from each participant when they were in other body postures, which included 54 10-second remounting sessions across three days. Hence, 864 minutes of data was collected in total for this user study. Note that the data P11 collected on Day 3 was lost due to the hardware issue so we asked the participant to come back again and redo the study on Day 3.
@@ 6.1. Study Results
In this subsection, we analyzed the performance of SonicID across multiple remounting sessions and days to simulate real-world wearing experience of smart glasses. The evaluation setup and results are visualized in Fig. 4.
@@ 6.1.1. Cross-session Performance
In the user study, we collected 36 sessions (6 minutes) of data from each participant while they were sitting still on each day. We first evaluated SonicID’s performance across different remounting sessions on the same day. Out of the 36 sessions we collected on Day 1, we considered the first 6 sessions (1 minute) of data as the practice sessions and dumped them while we evaluated the system. Among the remaining 30 sessions, we used the first 24 sessions (4 minutes) of data as positive samples to train the deep learning model described in Sec. 3.4. According to the data processing pipeline introduced in Sec. 3.3, we divided each 10-second session into multiple instances of 0.06 s and empirically picked top ten instances with the smallest movements by comparing the energy change within these instances. Ten 0.06-second instances from each session were actually used for training the model. The negative samples used to train the model came from a dataset we collected prior to the user study. To collect this dataset, we ran a pilot study of the same procedures as the formal user study with 21 people, including 7 researchers and 14 people from their networks. Among these people, 9 of them collected data on three days and the remaining 12 collected data on one day. Each person contributed 5 or 6 minutes of data of sitting still on each day. All these data in the dataset were used for training the base model in the pretraining stage and the individual models as negative samples. The same process was conducted to select the most static instances from all the sessions while training.
Figure 4. Evaluation Setup and Average Results of 40 Participants across Sessions and Days.
@@
After the model was properly trained for each participant in the user study, the remaining 6 sessions (1 minute) of data from this participant was adopted as positive samples to test the model. The negative samples for testing came from the other 39 participants in the user study. Each of these 39 participants contributed all sessions of data, excluding the first 6 sessions which were considered as practice sessions, to test the model. In this way, these "attackers" during evaluation were not seen by the model in the training process. Static instances were also selected from all sessions for testing. The model made binary classifications to authenticate users and computed the three metrics introduced in Sec. 3.5 to evaluate the results. We trained an individual model for each participant and the average TPR, FPR and BAC across 40 participants is 97.2% (std=9.0%), 4.6% (std=4.7%) and 96.3% (std=4.9%). The detailed metrics for every participant are shown in Figs. 7 to 9 respectively in Appendix A.
@@
The average FPR of 4.6% indicates that the system rejects false users with a success rate of over 95%. It demonstrates the ability of SonicID to protect the smart glasses from being attacked by bad actors. The average TPR of 97.2% means that the system authenticates the real users successfully with a rate of over 95%. It shows the ability of SonicID to recognize true users of the smart glasses. BAC gives a balanced evaluation of the two metrics above.
@@ 6.1.2. Cross-day Performance
The results above validated the performance of SonicID across different remounting sessions on the same day. Then we evaluated the performance across different days, which was meant to validate the stability of SonicID. On Day 2 and Day 3, we also collected 6 minutes of data from each participant when they were sitting still using the same procedures as Day 1. To be consistent with Day 1, we also dropped the first 1 minute of data from every participant. Then the static instances picked from the remaining data were used for testing the cross-day performance of SonicID. We did not retrain the models and directly used the models trained on Day 1. Each participant provided his/her own 5 minutes of data as positive samples to test the model for TPR on Day 2 and Day 3 separately. The negative samples still came from the other 39 participants in the user study to test FPR.
@@
The results showed that the average TPR, FPR and BAC for Day 2 across 40 participants are 81.3% (std=27.8%), 4.7% (std=4.5%) and 88.3% (std=14.5%) while those for Day 3 are 76.6% (std=29.2%), 4.0% (std=3.6%) and 86.3% (std=15.2%). The individual study results for each participant are displayed in Figs. 7 to 9 respectively in Appendix A. We ran a one-way repeated measures ANOVA test among TPR on three days for all 40 participants and identified a statistically significant difference ($F(2,117)=11.39,p=0.00005<0.05$). Specifically, we conducted a repeated measures t-test between TPR on Day 1 and Day 2, and we found a statistically significant difference ($t(39)=3.50,p=0.0012<0.05$). While the same repeated measures t-test was conducted between TPR on Day 1 and Day 3, a statistically significant difference was found ($t(39)=4.29,p=0.0001<0.05$). This indicates that TPR decreases on both Day 2 and Day 3, compared to that on Day 1. However, the average TPR is still over 75% even on Day 3. Under the cases when the real users fail to authenticate themselves into the smart glasses, they can adjust the glasses position or remount the device once, and they will likely log into the system successfully. The decrease of TPR can be attributed to the lack of positive training data from the user. One potential solution to improving TPR is to continuously collect the biomentric information of users’ face while they are using the device and update the model accordingly to increase the success rate of authentication. We explored this method in Sec. 6.4.
@@
Similarly, we also ran a one-way repeated measures ANOVA test among FPR on three days for all 40 participants and did not find a statistically significant difference ($F(2,117)=2.70,p=0.07>0.05$). This validates the stability of SonicID to protect the system from attackers across different days. In such authentication systems, it is usually more crucial to keep FPR lower than keeping TPR higher in case of a tradeoff since we generally do not want other people to have easy access to our personal devices.
@@
In Fig. 7 and Fig. 8, we can see that P32 has the lowest TPR on Day 2 and P30 has the lowest TPR on Day 3 while P4 has the highest FPR on Day 2 and P27 has the highest FPR on Day 3. By checking the videos of the participants recorded during the study for reference, we figured that P4, P27, P30 and P32 all have long hair and bangs which made it harder for them to put the glasses back to the same position every time after remounting. This negatively affected the performance of the system on them. In Sec. 7.4, we discussed several potential methods to improve the system performance and stability in the future.